# Agents Under Siege: Breaking Pragmatic Multi-Agent LLM Systems with

### v1

@@ Abstract
Most discussions about Large Language Model (LLM) safety have focused on single-agent settings but multi-agent LLM systems now create novel adversarial risks because their behavior depends on communication between agents and decentralized reasoning. In this work, we innovatively focus on attacking pragmatic systems that have constrains such as limited token bandwidth, latency between message delivery, and defense mechanisms. We design a permutation-invariant adversarial attack that optimizes prompt distribution across latency and bandwidth-constraint network topologies to bypass distributed safety mechanisms within the system. Formulating the attack path as a problem of maximum-flow minimum-cost, coupled with the novel Permutation-Invariant Evasion Loss (PIEL), we leverage graph-based optimization to maximize attack success rate while minimizing detection risk. Evaluating across models including Llama, Mistral, Gemma, DeepSeek and other variants on various datasets like JailBreakBench and AdversarialBench, our method outperforms conventional attacks by up to 7×7$\times$7 ×, exposing critical vulnerabilities in multi-agent systems. Moreover, we demonstrate that existing defenses, including variants of Llama-Guard and PromptGuard, fail to prohibit our attack, emphasizing the urgent need for multi-agent specific safety mechanisms.
@@
[…]
WARNING: This paper contains text that may be considered offensive.
Rana Muhammad Shahroz Khan1, Zhen Tan2, Sukwon Yun1, Charles Flemming3, Tianlong Chen1 1University of North Carolina at Chapel Hill, 2Arizona State University, 3Cisco
@@ 1 Introduction
To validate our method, we conduct extensive experiments across multiple architectures, including Llama-2-7B (Touvron et al., 2023). We benchmark our attack on a variety of datasets including JailbreakBench (Chao et al., 2024), demonstrating that our method achieves upto 7×7$\times$7 × the attack success rate compared with a vanilla prompt. Furthermore, we evaluate the effectiveness of existing safety mechanisms, including Llama-Guard (Inan et al., 2023), and show that they fail to defend against our attack in multi-agent settings. Lastly, we justify our settings and hyper-parameters with different ablation studies.
@@ Appendix B Experimental Settings
In this section we list down all the experimental settings, including datasets, architectures utilized, baselines and metrics. Compute: We utilize 8×8$\times$8 × Nvidia A6000s for all of our experiments.

### v2

@@ Abstract
Most discussions about Large Language Model (LLM) safety have focused on single-agent settings but multi-agent LLM systems now create novel adversarial risks because their behavior depends on communication between agents and decentralized reasoning. In this work, we innovatively focus on attacking pragmatic systems that have constrains such as limited token bandwidth, latency between message delivery, and defense mechanisms. We design a permutation-invariant adversarial attack that optimizes prompt distribution across latency and bandwidth-constraint network topologies to bypass distributed safety mechanisms within the system. Formulating the attack path as a problem of maximum-flow minimum-cost, coupled with the novel Permutation-Invariant Evasion Loss (PIEL), we leverage graph-based optimization to maximize attack success rate while minimizing detection risk. Evaluating across models including Llama, Mistral, Gemma, DeepSeek and other variants on various datasets like JailBreakBench and AdversarialBench, our method outperforms conventional attacks by up to $7\times$, exposing critical vulnerabilities in multi-agent systems. Moreover, we demonstrate that existing defenses, including variants of Llama-Guard and PromptGuard, fail to prohibit our attack, emphasizing the urgent need for multi-agent specific safety mechanisms.
@@
o  WARNING: This paper contains text that may be considered offensive.
Rana Muhammad Shahroz Khan1, Zhen Tan2, Sukwon Yun1, Charles Fleming3, Tianlong Chen1 1University of North Carolina at Chapel Hill, 2Arizona State University, 3Cisco
@@ 1 Introduction
To validate our method, we conduct extensive experiments across multiple architectures, including Llama-2-7B (Touvron et al., 2023). We benchmark our attack on a variety of datasets including JailbreakBench (Chao et al., 2024), demonstrating that our method achieves upto $7\times$ the attack success rate compared with a vanilla prompt. Furthermore, we evaluate the effectiveness of existing safety mechanisms, including Llama-Guard (Inan et al., 2023), and show that they fail to defend against our attack in multi-agent settings. Lastly, we justify our settings and hyper-parameters with different ablation studies.
@@ 6 Conclusion
Acknowledgment
This research was, in part, funded by the CISCO Faculty Award, UNC SDS Seed Grant and NetMind.AI. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing official policies, either expressed or implied of the funding organizations.
@@ Appendix B Experimental Settings
In this section we list down all the experimental settings, including datasets, architectures utilized, baselines and metrics. Compute: We utilize $8\times$ Nvidia A6000s for all of our experiments.